OpenStudy (anonymous):

A line passes through the point (1,6) and has a slope of -9 . Write an equation for this line

OpenStudy (anonymous):

please help!!

OpenStudy (mrnood):

Do you know the equation y=mx+b for all straight lines?

OpenStudy (anonymous):

I am to write the equation

OpenStudy (mrnood):

I fully understand that. The equation y=mx+b is a general equation for ALL straight lines. We need to know that before we can solve your question. Do you recognise that equation?

OpenStudy (anonymous):

no, i'm lost

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -9 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

ok, thx

OpenStudy (mrnood):

I'm not sure you truly understand that. You need to study some more about straight lines before trying this question (The formula I gave is basic to this subject so oyu need ot understand that before trying the answer. THIS page is a good starting place: http://www.mathsisfun.com/equation_of_line.html

OpenStudy (anonymous):

ok,thx